What is the standard form of 1(4+i) can you please give me the steps so I can learn how to do it
Answer by stanbon(75887)

Did you mean the trigonometric form of 4 + i?
---
If so:
r = sqrt(4^2+1^2) = sqrt(17)
theta = tan^1(1/4) = 14.04 degreees
-----
4+i = sqrt(17)[cos(14.04)+i*sin(14.04)]
=====
Cheers,
Stan H.